A handful of 21 nickels and dimes is worth $1.65. Find the number of each type of coin. Show your work.

Answer by TimothyLamb(4379)

5n + 10d = 165
n + d = 21
n = 21 - d
---
5n + 10d = 165
5(21 - d) + 10d = 165
105 - 5d + 10d = 165
5d = 60
---
d = 12
n = 9
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php